Problem 2 (60 POINTS)
(20 Points for each question)
Fig 1a shows a hollow tube with internal diameter of 230 mm and thickness of 30 mm, as shown in Fig 1b. The length of the tube is 1m.
Fig 1c shows a solid circular bar which also has the same length and the same weight as the hollow circular bar. Both the bars are made of same material (have same density or specific weight)
Fig 1a
Fig 1b Fig 1c
a) Calculate the radius of the solid circular bar.
b) Calculate the Polar moment of Inertia for both the bars.
c) Mention which one has a bigger value of Polar Moment of Inertia (J). Explain in details the significance
of having a bigger value of J considering both the bars are only under torsional load. Your explanation may be supported by formulas and illustrations.
Answer #1

D = 176.635 mm Solution page so Solid bar Hollow tube 2.1) 30mm Length Do Di +2+ = 230 + 2x30 = 240 mm bar. given weight of h1%, (176.839) m on bar we T 2) TE TJ which means hollow Shaft can dolution page no 2 For solid bar Is = 1/2D = Jş = 9.5566 7Solution page no: 3 general cax Do Di tone7 0 @ consider hollow Shaft outer diameter la and diameter Di re Di De Similarly co